OpenStudy (anonymous):

Simplify the following expression. a + 2c - 5b - c + a - b A.)2a - 6b + 2c B.)2a - 6b + c C.) a - 6b + c D.) a - 5b + 2c

OpenStudy (pfenn1):

a + 2c - 5b - c + a - b = a + a - 5b - b + 2c - c =?

OpenStudy (anonymous):

2a-6b+c